Tags:
create new tag
view all tags

Shift Log 20101207

0520

Changed kicker timing back to the original after FONT shift. The number is 0591. The damping ring was also put back into its original state, since the tune had been changed. The values are set to

H: 0.180547 V: 0.561389

0525

Looking at orbit in EXT and FF

0527

Improving transmission by tweaking injector with multi knobs. The following were adjusted in a loop:

Laser -> Gun -> L0

whilst observing the DR charge monitor.

0530

Tuning the orbit in EXT by changing the steering magnets.

0535

Tuning the gun correctors (magnets beginning with ST).

0540

Checked background in detector. ~300-350 mV, corresponding to ~10000 ADC counts with no attenuation. Opened the EDM display in the control room after new PVs were added.

0550

Checking that QM16FF mathches the original optics calculated by Woodley.

0557

Adjusted angle at LW1X, but downstream orbit hitting aperture (could be seen in EXT LOSS-MON display).

0623

Since we are saturating the ADC all the time, some attenuation (16 dB) was put in.

0635

We now have orbit and "golden" orbit from last shift in the EDM display, and are using it to tune the orbits. Trying to adjust the positions of magnets to correct the orbit.

QF7FF moved from 0 to +1500 microns. QF3FF moved from 0 to 450 microns.

0810

Tuning the orbit has made little difference to the Cherenkov detector background, but reducing QM16FF to ~50% brought the detector counts back to ~10000. Moving QM16FF by a large displacement (~200 microns) did not change the background, though, so is not believed to be steer-related.

Vertical: QS1X and QS2X. Horizontal: QF1X, QF6X. These are the knobs that change the dispersion function. We are changing these with QM16FF at different settings to see if dispersion 'talks' to QM16FF.

0820

Some magnets were set to the wrong settings (not the settings calculated in MAD). QM16FF should have been 15.766 A. QF19X was different by a few amperes. A couple of other magnets have the wrong settings (see saved files). The CORRECT magnets settings were saved in a new set file: set10dec07_0826.dat (NOTE QM16FF was not saved in the file due to a GUI error. It's setting was 15.766 A). Now we have also have "good" (~10000 ADC counts, no attenuation) background, and QM16FF has the correct setting.

Wire scans

Axis #sigma [#mum] <#sigma> [mm]
MW1X y 37.4 +/- 8.3 53.162 +/- 0.008
x 146.5 +/- 14.8 20.764 +/- 0.012
MW2X y 11.7 +/- 0.7 62.158 +/- 0.001
x 68.2 +/- 4 34.132 +/- 0.003

0900

Another saved set optics file. (Same optics): set10dec07_0857.dat. Again, NOTE QM16FF will not be shown correctly in this file. Its setting was 15.766 A.

Tuning up charge again.

0930

Charge ready, now setting timing for collisions using knife edge and APD (usual procedure).

12:00

Phase scan LTD2 : 813, PV : 5.8 lwPhaseScan_20101207_121228.dat

12:30

CH : 100 CV: [320,380,30,5] lwIpMoverScan_20101207_122154.dat

12:40

Added 30 cm of cable to main laser TD2 start

CH : 100 um CV : [280,420,20,5] lwIpMoverScan_20101207_124537.dat 28 um
CH : [-400,600,20,5] CV : 349 lwIpMoverScan_20101207_125604.dat Not complete range
CH : [-600,900,30,5] CV : 349 lwIpMoverScan_20101207_125604.dat
CH : [-600,1200,50,5] CV : 349 lwIpMoverScan_20101207_130753.dat
CH : [-600,1200,50,5] CV : lwIpMoverScan_20101207_133009.dat
CV : [280,420,70,5] CV : 348 lwIpMoverScan_20101207_134024.dat
PH : [4,7,30,5] CV : 5.61 lwPhaseScan_20101207_135125.dat

Signal vs CLaser Energy
50 pulse logs

0.00008 s lwLog_20101207_141414.dat
0.00009 s lwLog_20101207_141539.dat
0.00010 s lwLog_20101207_141624.dat
0.00011 s lwLog_20101207_141709.dat
0.00012 s lwLog_20101207_141803.dat
0.00013 s lwLog_20101207_141843.dat
0.00014 s lwLog_20101207_141925.dat
0.00015 s lwLog_20101207_142011.dat
0.00016 s lwLog_20101207_142111.dat
0.00017 s lwLog_20101207_142235.dat
0.00018 s lwLog_20101207_142344.dat
0.00019 s lwLog_20101207_142447.dat

100 pulses - didn't stop laser, but around 300uJ so negligible. 0.00150 s lwLog_20101207_142625.dat

14:30 - 15:00

Tried to do signal vs ATF charge (via gun laser intensity) but kept losing and regaining collisions. REF1:pha - phase of reference 1 bpm phase shows a jump of around 0.3rad. Added to log for correlation.

Going to proceed with laser waist scanning ie vertical horizontal and vertical etc

15:00 laser waist profiling

15:30 Access to restart computer. Laser energy VI crashes computer when scanning the chamber.

Laser energy around 100mJ

Uncertainties from fitting.

CV : [280,420,70,5] CH 750 lwIpMoverScan_20101207_160049.dat Unclear from data. Signal to background and noise too low.
CV : [280,420,70,5] CH 650 lwIpMoverScan_20101207_160851.dat sigma = 16.73 +-, centre = 345.78 +- 0.17
CV : [280,420,70,5] CH 550 lwIpMoverScan_20101207_155107.dat sigma = 10.55 +- 0.7, centre = 338.9 +- 0.07
CV : [280,420,70,5] CH 450 lwIpMoverScan_20101207_161838.dat sigma = 10.55 +- 0.7, centre = 338.9 +- 0.07
CV : [280,420,70,5] CH 350 lwIpMoverScan_20101207_153943.dat sigma = 13.7 +- 0.07, centre = 340.78 +- 0.06
CV : [280,420,70,5] CH 250 lwIpMoverScan_20101207_163248.dat sigma = 18.59 +- 0.1, centre = 337.6 +- 0.07
CV : [280,420,70,5] CH 150 lwIpMoverScan_20101207_154534.dat sigma = 17.62 +- 0.1, centre = 334.9 +- 0.08
Plot doesn't show clear minima. Why is this? Jitter? Drift? Odd shaped beams? Horizontal scan to find out.
CH : [0,1000,10,5] CV 342 lwIpMoverScan_20101207_164125.dat Not clear - peak around 200 - extending range
CH : [0,1000,10,5] CV 342 lwIpMoverScan_20101207_165055.dat

Signal vs Charge - previous try wasn't good as timing kept us slipping out of collisions. 500 Pulse log whilst continuously decreasing gun laser intensity down form 20% to 8% (non-linear).

lwLog_20101207_165904.dat

Signal vs Energy. Lowest possible signal seen at 20mJ. 100 pulse log taken whilst switching laser on and off.

lwLog_20101207_171312.dat

17:37 GW MW Load matched optics

  • Loaded optics, no change in background collisions still stable
  • standardize magnets

18:22 First scan

../dat/ana/xyfitplot_lwIpMoverScan_20101207_182022.pdf

20:00

Glen White (GW) & Mark Woodley (MW) did a dispersion correction. New dispersion at our IP is x: 1.1 +- 0.5mm and y: 0.06 +- 0.09mm.

First vertical scan
CV : [280,420,35,5] CH 250 ../dat/raw/lwIpMoverScan_20101207_200058.dat
sigma = 9.78 +- 4.51, centre = 320.6 +- 4.37
Second vertical scan
CV : [230,410,35,5] CH 250 ../dat/raw/lwIpMoverScan_20101207_200755.dat
sigma = 8.58 +- 4.63, centre = 322.2

GW looked at OTR#0 and MW adjusted QK1X from 0A to -15.5A.

CV : [250,370,120,5] CH 250 ../dat/raw/lwIpMoverScan_20101207_203614.dat
sigma = 8.54 +- 0.04, centre = 319.13 +- 0.04
Saturated though at the peak of the signal. Additionally, it was slightly asymmetrical so the Gaussian fit didn't reach the top. Increase attenuation to ADC from 4dB to 8dB.

CV : [270,370,100,5] CH 250 ../dat/raw/lwIpMoverScan_20101207_205242.dat
sigma = 7.89 +- 0.04, centre = 318.92 +- 0.04

CH : [0,500,25,5] CH 250 ../dat/raw/lwIpMoverScan_20101207_210103.dat
sigma = 84.5 +- 3.4, centre = 255 +- 2

CV : [270,370,100,5] CH 255 ../dat/raw/lwIpMoverScan_20101207_210903.dat
sigma = 7.96 +- 0.05, centre = 316.11 +- 0.05

Edit | Attach | Watch | Print version | History: r9 < r8 < r7 < r6 < r5 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r9 - 13 Apr 2012 - LauraCorner

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback